OpenStudy (potatoofhistory):

WILL GIVE MEDAL AND FAN PLEASE HELP!!!!!! A bucket that contains several inches of water is left outside on a rainy day. The height of the water in the bucket that day is given by the function equation H = 4 + 0.5t, where H is the height of the water and t is the time in hours the bucket sits out in the rain. What is the meaning of the vertical intercept in this equation? A. The bucket originally contained 0.5 in. of water. B. The bucket originally contained 4 in. of water. C. The bucket fills at a rate of 0.5 in./h. D. The bucket fills at a rate of 4 in./h.

OpenStudy (anonymous):

B

OpenStudy (potatoofhistory):

Are You Sure?

OpenStudy (potatoofhistory):

I can't fail this test.

OpenStudy (anonymous):

an equation is y = mx + b and the Y intercept is B so if we reform the equation: H = 4 + 0.5t into H = 0.5t + 4 then 4 is the Y intercept

OpenStudy (anonymous):

and the 4 represents the original height of the water in the buchet

OpenStudy (potatoofhistory):

Can you help me with 2 more questions?

OpenStudy (anonymous):

sure

OpenStudy (potatoofhistory):

Marcus is eating a meal that is made up of an entrée that is 45 calories per ounce and a side dish that is 30 calories per ounce. He wants his meal to total 510 calories. Let e represent the number of ounces of the entrée that he eats, and let s represent the number of ounces of the side dish that he eats. Which function relates the number of ounces of the side dish that he eats to the number of ounces of the entrée that he eats? A. S= 2/3e -17 B. S=3/2e - 11 and 1/3 C. S= 2/3e + 11 and 1/3 D. S= 3/2e + 17

OpenStudy (potatoofhistory):

when i do 2/3 i mean 2 over 3

OpenStudy (anonymous):

what is entr������������������e?

OpenStudy (potatoofhistory):

?

OpenStudy (potatoofhistory):

Marcus is eating a meal that is made up of an entrée that is 45 calories per ounce and a side dish that is 30 calories per ounce. He wants his meal to total 510 calories. Let e represent the number of ounces of the entrée that he eats, and let s represent the number of ounces of the side dish that he eats. Which function relates the number of ounces of the side dish that he eats to the number of ounces of the entrée that he eats? idk if you see the<?> but i see it just fine.

OpenStudy (anonymous):

woah i didn't do anything what happened?

OpenStudy (potatoofhistory):

ugh hold on i know whats wrong

OpenStudy (potatoofhistory):

Marcus is eating a meal that is made up of an entree that is 45 calories per ounce and a side dish that is 30 calories per ounce. He wants his meal to total 510 calories. Let e represent the number of ounces of the entree that he eats, and let s represent the number of ounces of the side dish that he eats. Which function relates the number of ounces of the side dish that he eats to the number of ounces of the entree that he eats?

OpenStudy (anonymous):

I'm really sorry but i don't know how to solve that one... I can't make sense of it

OpenStudy (potatoofhistory):

it was the weird e's in entree

OpenStudy (potatoofhistory):

there is the medal for the first. thanks for the help anyways.

OpenStudy (anonymous):

thanks